OpenStudy (anonymous):

The sun roof in Troy’s house is shaped like a rectangle. He increases the length of the sides of the sun roof to four times the existing ones. How will this change affect the perimeter of the sun roof?

OpenStudy (anonymous):

P=2b+2h if b and h are multiplied by 4, our new perimeter S is: S=2(4b)+2(4h)=4(2b+2h)=4P So our new perimeter will be 4 times as big as the old one.

OpenStudy (anonymous):

Thankyou for the explanation

OpenStudy (anonymous):

np, hope it makes sense